What is the perimeter in meters of a rectangular board with a length of 47cm and a width of 1.6m?
Answer by oscargut(2103)

P = 2(L+W) = 2(0.47+1.6) = 4.14 m
Answer: 4.14 m
If you need more help with similar questions you can contact me now at:
mthman@gmail.com
Thanks